OpenStudy (anonymous):

Graph the function. y = |cot x/4|

geerky42 (geerky42):

\[y = \left| \cot\left( \dfrac{x}{4} \right) \right|\] http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=Graph%20the%20function.%20y%20%3D%20%7Ccot%20x%2F4%7C

OpenStudy (anonymous):

x can take values from 0 to infinity and u calculate ctg fron those values Cotangent(2.5 rad) = -1.338648 Cotangent(2.25 rad) = -0.807345 Cotangent(2 rad) = -0.457658 Cotangent(1.75 rad) = -0.181147 Cotangent(1.5 rad) = 0.070915 Cotangent(1.25 rad) = 0.332273 Cotangent(1 rad) = 0.642093 Cotangent(0.75 rad) = 1.073426 Cotangent(0.5 rad) = 1.830488 Cotangent(0.25 rad) = 3.916317 0.25 means x =1 so ctg of 1/4 and so on http://www.endmemo.com/trigonometry/cotangent.php

geerky42 (geerky42):

It is the parent function (cot x).

OpenStudy (anonymous):

cot is a function, y is the line that has as value a function, in this example cot and cot has parameters as x or x over 4 or anything that contains a variable

OpenStudy (anonymous):

ur answer is good geerky42, sry for the misleading information i now have seen my mistake

OpenStudy (anonymous):

These are my options:

OpenStudy (dumbcow):

its graph A because range must be greater than 0 and there is asymptote at x=0

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw